OpenStudy (brycedowling2002):

Simplify. 3[5 + (10 – 3)^2]

OpenStudy (mathstudent55):

Use the order of operations, PEMDAS. Start working inside the parentheses. What is 10 - 3?

OpenStudy (brycedowling2002):

7

OpenStudy (mathstudent55):

Good. Now that is inside parentheses and being raised to the 2nd power. What is 7^2?

OpenStudy (brycedowling2002):

49

OpenStudy (brycedowling2002):

hello

OpenStudy (brycedowling2002):

im stuck

OpenStudy (mathstudent55):

Good. So far you're doing well. 3[5 + (10 – 3)^2] = = 3[5 + (7)^2] = 3[5 + 49] We are here now. The next step is to work inside the brackets. What is 5 + 49?

OpenStudy (brycedowling2002):

54

OpenStudy (mathstudent55):

Great. Only one step left. = 3[5 + 49] = 3[54] Now multiply 3 by 54

OpenStudy (brycedowling2002):

162

OpenStudy (brycedowling2002):

thats my answer

OpenStudy (mathstudent55):

Correct.

OpenStudy (brycedowling2002):

thank u